OpenStudy (anonymous):

an engineer is designing the runway of an airport.Of the planes which will use the airport, the lowest acceleration rate is 3m/s^2 and the lowest take-off speed is 65 m/s.assuming these minimum parameters are for the same plane, what is the minimum allowed length of the runway?

OpenStudy (anonymous):

How far will it take this plane to reach 65 m/s as final V with original Vo as 0m/s and acceleration at 3m/s^2? Don't worry, that's what you have me for. V=Vo+a*t since Vo is 0m/s, V/a=t Now, we know the equation D=(V+Vo)*t/2, so plug in 0 for Vo and V/a for t to get.. D=(V)(V/a)/2=65^2/(3*2)=704m